OpenStudy (anonymous):

need to find if converges or diverges using ratio test: sum of all (ln(n))/(n) when n=1 to inf *not sure what cancels once RT is set up

OpenStudy (anonymous):

any idea?

OpenStudy (anonymous):

...why the ratio test? painful.

OpenStudy (anonymous):

The ratio I get using the ratio test is 1. It wouldn't be a conclusive test.

OpenStudy (anonymous):

If you want me to show you that, let me know.

OpenStudy (anonymous):

it's not that painful loki all you've got to do is put : an+1 / an , I find it rather simple lol ^_^

OpenStudy (anonymous):

Yeah, I know ;)

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

\[|\frac{\ln (n+1))}{n+1}/\frac{\ln(n)}{n}|=\frac{\ln(n+1)}{\ln(n)}\frac{n}{n+1}\]This is an indeterminate form (infty/infty), so you can use L'Hopital's rule:\[\lim_{n->\infty}\frac{\ln (n+1)}{\ln(n)}\frac{n}{n+1}=\lim_{n->\infty}\frac{\ln (n+1)}{\ln (n)}\lim_{n->\infty}\frac{n}{n+1}\]\[=\lim_{n->\infty}\frac{\frac{1}{n+1}}{\frac{1}{n}}\lim_{n->\infty}\frac{1}{1}=\lim_{n->\infty}\frac{n}{n+1}=\lim_{n->\infty}\frac{1}{1+\frac{1}{n}}=1\]